Grade 10General Physics

what is Stefan - Boltzmann’s Law.it explain

HARINATH
10 Years agoGrade 10

11 Answers

mahendra
ApprovedApproved Tutor Answer10 Years ago
a) If a black body at absolute temperature T is surrounded by an enclosure at absolutetemperature To , then the rate of loss of heat energy by radiation per unit area is given by.E = σ (T 4 − T0 4 )b) For any hot body, E = σ Ae(T 4 − T0 4 )Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
venky
10 Years ago
he StefanBoltzmann law, also known as Stefan's law, describes the power ... Specifically, the StefanBoltzmann law states that the total energy radiated per ..
mahendra
10 Years ago
a) If a black body at absolute temperature T is surrounded by an enclosure at absolute
temperature To , then the rate of loss of heat energy by radiation per unit area is given by.
E = σ (T 4 − T0 4 )
b) For any hot body, E = σ Ae(T 4 − T0 4 )
Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
 
venky
10 Years ago
rey harinath approve chey
salvathshaik
10 Years ago
 If a black body at absolute temperature T is surrounded by an enclosure at absolutetemperature To , then the rate of loss of heat energy by radiation per unit area is given by.E = σ (T 4 − T0 4 )b) For any hot body, E = σ Ae(T 4 − T0 4 )Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
mahendra
10 Years ago
a) If a black body at absolute temperature T is surrounded by an enclosure at absolutetemperature To , then the rate of loss of heat energy by radiation per unit area is given by.E = σ (T 4 − T0 4 )b) For any hot body, E = σ Ae(T 4 − T0 4 )Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
salvathshaik
10 Years ago
 If a black body at absolute temperature T is surrounded by an enclosure at absolutetemperature To , then the rate of loss of heat energy by radiation per unit area is given by.E = σ (T 4 − T0 4 )b) For any hot body, E = σ Ae(T 4 − T0 4 )Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
mahendra
10 Years ago
a) If a black body at absolute temperature T is surrounded by an enclosure at absolutetemperature To , then the rate of loss of heat energy by radiation per unit area is given by.E = σ (T 4 − T0 4 )b) For any hot body, E = σ Ae(T 4 − T0 4 )Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
pa1
10 Years ago
 
a) If a black body at absolute temperature T is surrounded by an enclosure at absolute
temperature To , then the rate of loss of heat energy by radiation per unit area is given by.
E = σ (T 4 − T0 4 )
b) For any hot body, E = σ Ae(T 4 − T0 4 )
Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body
 
sathish
10 Years ago
a) If a black body at absolute temperature T is surrounded by an enclosure at absolute temperature To , then the rate of loss of heat energy by radiation per unit area is given by. E = σ (T 4 − T0 4...
Surendra
10 Years ago
a) If a black body at absolute temperature T is surrounded by an enclosure at absolutetemperature To , then the rate of loss of heat energy by radiation per unit area is given by.E = σ (T 4 − T0 4 )b) For any hot body, E = σ Ae(T 4 − T0 4 )Where ‘e’ is the emissive power and ‘A’ is the area of cross-section of the hot body